6. Find the exact value of ,* f(x) dx, if the graph of f(x) is given below. 6 5 3 3 2 1 0 2 3 4 5 6 7 8 9
Answer #1

Ans) now we see

\int_{I}^{8}f'(x){dx}=\left | f(x) \right |_{1}^{8}=\left ( f(8)-f(1) \right )

as given by graph f(8) = 5

and f(1) = 1

therefore we have

\int_{I}^{8}f'(x){dx}=\left ( f(8)-f(1) \right ) = 5 - 1 = 4

thus the value of integration of f'(x) between x=1 to x =8 is 4